OpenStudy (anonymous):

use the discriminant to find the number and kind of solutions for the equation 2x2 + 7x = 10

OpenStudy (anonymous):

The answer is this : 4+7x=10 7x=10-4 7x=6 x=6/7 Did you want it ?!:)

OpenStudy (anonymous):

i wrote that equation wrong.. it is \[2x ^{2} + 7x = 10\]

OpenStudy (anonymous):

Is this your question ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok ! We have : 2x^2+7x-10=0

OpenStudy (anonymous):

It s a 2th equetion ! You should find X in this way but remamber we have -x and +x .!

OpenStudy (anonymous):

Firs find \[Deltab\]: To find it can use this formula : Delta = b^2-4ac 2=a -10=c 7=b Then to find x : Understand ?

OpenStudy (anonymous):

@annalizz12

OpenStudy (anonymous):

D=b^2-4ac=7^2-4*2*(-10)=49+80=139 If D=0 roots are real and equal if D is positive and a perfect square,then roots are real rational and different. If D is positive and not a perfect square roots are real ,irrational and different. If D is negative, then roots are not real,they are imaginary.

OpenStudy (anonymous):

hey tthe answer is that u have to ask ur theacher just kidding lol:)

OpenStudy (anonymous):

correction 49+80=129

OpenStudy (anonymous):

ok ok ok ok ok ok ok ok ok

OpenStudy (anonymous):

you can see out of the four options which matches.with your question.

OpenStudy (anonymous):

yes ic